Reports until 09:36, Tuesday 04 June 2024
GQuEST Output Filter Cavities
Torrey Cullen - posted 09:36, Tuesday 04 June 2024 (11624)
Frequency shifted 775 light in filter cavity

The twice shifted light via the AOM has been aligned into the fiber that brings the light to the filter cavity. It looks like alignment is slightly worse after messing with the fiber, alignment touch up is needed. There is also a ghost beam from somewhere now that should be diagnosed. 775 light is now ready though.

Just so people know, there is a constant in frequency sine wave on the error signal of the 775 reflected light. After talking with Lee, he suggested it being a beat note of the EOMs. I turned off the 1550 EOM and the beatnote in the error signal disappears. Something to keep in mind going forward.